# Multi-Lattice Sampling of Quantum Field Theories via Neural Operator-based Flows

### 13 Jan 2024 v2

@@ 1 Introduction
Let $S$ an action on some lattice (as an approximation of the continuous action $\mathcal{S}$). Albergo et al. [1] propose to sample from the quantum field theory defined by $S$ by using a normalizing flow to parametrise a density $q_{\theta}$ on the lattice and optimize its parameters $\theta$ until $q_{\theta}$ is a good approximation of $Z^{-1}e^{-S}$.
Operator learning promotes the viewpoint that the lattice/mesh is merely a computational tool, and model should capture the underlying continuous physics. Kovachki et al. [2] term this property of models discretization invariance.  [^1] Discretization invariance means that the neural operator evaluated on finer and finer discretizations approximates the continuous operator. Thus, strictly speaking, it is not a requirement of invariance rather that of convergence.
In this work we apply the same idea to the task of sampling from lattice quantum field theories, motivated by the fact that lattice field theories also emerge as the discretization of continuous field theories. Suppose now that the field theory is defined on some domain $D$. Once a lattice, as a discretization of $D$, is chosen, one can construct a continuous normalizing flow [3] by a time-dependent vector field $V_{t}$ that parametrizes the direction along which probability mass moves. Generalizing this idea, we propose to parametrize a time-dependent operator $\mathcal{V}_{t}$ from the space of functions on $D$ to itself that defines the direction in which functional probability mass moves. Such an operator can then be used to map the functional distributions $[\mathcal{D}\phi(x)]\mathcal{Z}_{0}^{-1}e^{-S_{0}[\phi(x)]},[\mathcal{D}\phi(x)]\mathcal{Z}_{1}^{-1}e^{-S_{1}[\phi(x)]}$ to one another. Computationally the operator $\mathcal{V}_{t}$ can only be accessed by a choice of a lattice which induces a vector field $V_{t}$ as the discretization of $\mathcal{V}_{t}$. We then train this vector field to be a diffeomorphism between the discretized free and target theories, $[d\phi]Z_{0}^{-1}e^{-S_{0}[\phi]}$ and $[d\phi]Z_{1}^{-1}e^{-S_{1}[\phi]}$. The upside of using a operator-based flow will be that a single model can be used to operate on multiple discretizations of the same underlying continuous system. Figure 1 provides a schematic overview of the objects and their relation in this paragraph.

### 17 Jan 2024 v3

@@ 1 Introduction
Consider $\mathcal{S}$ as an action characterizing a quantum field theory, with $S$ representing its discretization to a lattice. Albergo et al. [1] suggest a method for sampling from the lattice quantum field theory described by $S$. This involves a normalizing flow parameterizing a probability density $q_{\theta}(\phi)$ of discrete fields over the lattice, and optimizing the parameters $\theta$ until $q_{\theta}(\phi)$ closely approximates $\frac{e^{-S[\phi]}}{Z}$, where $Z=\int[d\phi]e^{-S[\phi]}$ is the normalizing constant.
Operator learning promotes the viewpoint that the lattice/mesh is merely a computational tool, and model should capture the underlying continuous physics. Kovachki et al. [2] term this property of models discretization invariance.  [^1] Discretization invariance means that the neural operator evaluated on finer and finer discretizations approximates the continuous operator. Thus, strictly speaking, it is not a requirement of invariance rather that of convergence. In this work, we apply the same idea to the task of sampling from lattice quantum field theories, motivated by the fact that lattice field theories also emerge as the discretization of continuous field theories.
Suppose now that the field theory is defined on some domain $D$. Once a lattice, as a discretization of $D$, is chosen, one can construct a continuous normalizing flow [3] as a time-dependent vector field $V_{t}$ that parametrizes the direction along which probability mass moves. Generalizing this idea, we propose to parametrize a time-dependent operator $\mathcal{V}_{t}$ from the space of functions on $D$ to itself that defines the direction in which functional probability mass moves. Such an operator can then be used to map the functional distributions $[\mathcal{D}\phi(x)]\mathcal{Z}_{0}^{-1}e^{-S_{0}[\phi(x)]},[\mathcal{D}\phi(x)]\mathcal{Z}_{1}^{-1}e^{-S_{1}[\phi(x)]}$ to one another. Computationally the operator $\mathcal{V}_{t}$ can only be accessed by a choice of a lattice which induces a vector field $V_{t}$ as the discretization of $\mathcal{V}_{t}$. We then train this vector field to be a diffeomorphism between the discretized free and target theories, $[d\phi]Z_{0}^{-1}e^{-S_{0}[\phi]}$ and $[d\phi]Z_{1}^{-1}e^{-S_{1}[\phi]}$. The upside of using a operator-based flow will be that a single model can be used to operate on multiple discretizations of the same underlying continuous system. Figure 1 provides a schematic overview of the objects and their relation in this paragraph.